### v1

# Optimization-based control by interconnection of nonlinear port-Hamiltonian systems

Abstract.
In this paper, we formulate an optimization-based control-by-interconnection approach to the stabilization problem of nonlinear port-Hamiltonian systems. Motivated by model predictive control, the feedback is defined as an initial part of a suboptimal solution of a finite horizon optimal control problem. To this end, we write the optimization method given by a primal-dual gradient dynamics arising from a possibly control-constrained optimal control problem as a port-Hamiltonian system. Then, using the port-Hamiltonian structure of the plant, we show that the MPC-type feedback law is indeed a structure-preserving interconnection of two port-Hamiltonian systems. We prove that, under an observability assumption, the interconnected system asymptotically stabilizes the plant dynamics. We illustrate the theoretical results by means of a numerical example.
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 531152215 – CRC 1701.
Keywords: Port-Hamiltonian systems, control by interconnection, suboptimal Model Predictive Control, monotone operators, nonlinear control, optimal control
1. Introduction
Many real-world phenomena and industrial applications can be modeled as nonlinear control systems. A central and ongoing challenge in this setting is to design control laws that simultaneously stabilize the system, optimize a prescribed performance criterion, and satisfy given constraints. In the presence of non-quadratic cost functions, nonlinear dynamics and state or control constraints, standard approaches such as the state-feedback law derived from the solution of the algebraic Riccati equation are no longer applicable. To address these limitations, nonlinear optimization-based control schemes such as model predictive control (MPC) have emerged as well-suited alternatives. In MPC, the current state of the system is measured (or estimated), a finite-horizon optimal control problem (OCP) subject to the system dynamics and constraints is solved, and a first portion of the resulting optimal control sequence is applied to the plant, see Figure 1 for an illustration. Under suitable assumptions, an iterative application of this procedure leads to a stabilizing controller for the plant, see the textbooks [13, 19].
$\dot{x}_{p}(t)=f_{p}(x_{p}(t),u_{p}(t))$ $\min$ $\qquad J(x,u)$ $\mathrm{s.t.}\ \dot{x}(\tau)$ $=f(x(\tau),u(\tau)),\ x(0)=x_{0}$ $(x,u)$ $\in\mathbb{X}\times\mathbb{U}$ $x_{0}=x_{p}(t)$$u_{p}(t)=u^{\star}(0)$
Figure 1. Illustration of Model Predictive Control
However, for small sampling times and in view of increasing complexity of real-world applications such as problems governed by partial differential equations, it can be challenging to compute the optimal solution in a short time. As a consequence, there is need for MPC approaches that utilize suboptimal solutions as feedback, obtained e.g. by an iterative optimization algorithm. In these methods, the solution of the optimal control problem on the bottom right of Figure 1 is replaced by an inexact solution given, e.g., by the truncation of an iterative optimization method. In recent decades, various suboptimal MPC schemes for finite-dimensional nonlinear control systems have been proposed and analyzed [8, 29, 11, 20].
In this work, we introduce a suboptimal MPC-like control scheme tailored to port-Hamiltonian systems (pHs). PHs [15] have emerged as a powerful modeling framework for physical systems and constitute a sub-class of dissipative systems [4, 28]. They provide a unified energy-based representation that naturally encodes passivity, interconnection structure, and dissipation [24, 14, 16]. Moreover, the port-Hamiltonian formalism offers powerful guarantees on stability and robustness under structured interconnections, making it particularly attractive for control design, e.g., via control by interconnection (CbI; [17]). Control by interconnection (CbI) for pHs is a control paradigm in which a controller, modeled itself as a passive system, is interconnected with the plant in a power-preserving manner. The resulting closed-loop system inherits passivity, with its total energy given by the sum of the individual storage functions, enabling energy shaping and, in particular, stabilization. However, achieving stabilization for port-Hamiltonian systems with input constraints remains a nontrivial task, as classical passivity-based methods typically do not address inequality constraints.
To address this issue, we propose a control by interconnection approach in the spirit of suboptimal MPC by formulating a continuous-time primal-dual gradient method for nonlinear control-constrained optimal control problems as a port-Hamiltonian system. To this end, we interconnect the dynamics of gradient-type optimization algorithm via an MPC-type feedback with a port-Hamiltonian plant in a structure-preserving-way, leading to a dissipative even asymptotically stable closed-loop system depicted in Figure 2. We provide stability results for primal-dual gradient dynamics and prove convergence of the interconnected system in function space. This is particularly central as the convergence does not rely on any particular sampling or discretization scheme, as long as it preserves the port-Hamiltonian structure. Last, the proposed analysis in function space paves the way towards suboptimal MPC for partial differential equations, a topic that has been largely unexplored so far. In this context we mention related works. Finite-dimensional saddle-point dynamics occurring in primal-dual gradient methods have been analyzed in [7, 6] and the port-Hamiltonian structure has been observed in [22] for stationary finite-dimensional problems. In [18, 26], the authors suggest an MPC-type control-by interconnection of port-Hamiltonian systems for discrete-time finite-dimensional optimal control problems.
$\dot{x}_{p}(t)$ $=-M_{p}(x(t))+B_{p}u_{p}(t)$ $y_{p}(t)$ $=B_{p}^{\top}x_{p}(t)$ $\dot{w}(t)$ $=-M_{\mathrm{opt}}(w(t))+\mathcal{B}u_{\mathrm{opt}}(t)$ $y_{\mathrm{opt}}(t)$ $=\mathcal{B}^{*}w(t)$ $u_{\mathrm{opt}}(t)=By_{p}(t)$$u_{p}(t)=-B^{\top}\!y_{\mathrm{opt}}(t)$
Figure 2. Proposed MPC-type control-by-interconnection scheme. The subscript $\mathrm{p}$ refers to the plant to be controlled where $\mathrm{opt}$ denotes the dynamics of state, control and adjoint in the optimization scheme.
Our main contribution, situated at the interface of control-by-interconnection and structure-exploiting optimal control for pHs, and the structure of this work are as follows. First, in Section 2, we introduce a class of infinite-dimensional port-Hamiltonian systems subject to set-valued nonlinear operators that is suitable to model primal-dual gradient dynamics for continuous-time control-constrained optimal control problems. The second contribution is centered around formulation (Section 3) and analysis (Section 4) of the primal-dual gradient dynamics for continuous-time control-constrained problems. In Theorem 4.7, we establish weak convergence of the primal–dual gradient dynamics to the unique optimal point of the underlying optimization problem. The proof is based on an infinite-dimensional LaSalle invariance principle leveraging the convexity of the problem and a detectability-like assumption of the cost function. The corresponding result for control-constrained problems is given in Theorem 4.8. Third, we prove stability of a suboptimal MPC-type control-by-interconnection method for unconstrained stabilization of pHs in Section 5, namely in Theorems 5.2 and 5.3 based on either coervicity or system-theoretic properties such as detectability and controllability using again a LaSalle-type argument. An extension to control-constrained stabilization is provided in Theorem 5.6 under an observability condition. We conclude our work by means of a numerical experiment in Section 6. The proofs of two theorems related to our LaSalle-type result are moved to Appendix 8.1. Some important notions and results from convex analysis are recalled in Appendix 8.2.
2. M-monotone pH systems
In this section, we introduce a novel class of monotone port-Hamiltonian systems in Hilbert spaces. This class will be used to describe both, the optimizer dynamics of the gradient-type algorithm in function space as well as the plant (see Figure 2). After introducing the model class, we provide solvability results, analyze stability and passivity properties and last investigate structure-preserving interconnections. The latter will lay the foundation for our suggested suboptimal MPC-type control scheme defined in Section 5.
Monotone pHs have been examined in [5] for the finite-dimensional case and in [10] for infinite-dimensional unconstrained optimal control problems. Here, we study particularly the case of constrained optimal control problems for which the optimality system may be written in terms of multi-valued operators as shown later in Section 3.
We briefly introduce some required notation. Let $(X,\langle\cdot,\cdot\rangle)$ be a real Hilbert space and $M\subset X\times X$. Using the same symbol, we associate with this set a set-valued operator $M:X\rightrightarrows X$ where
$M(x)=\left\{y\in X\,\middle|\,(x,y)\in M\right\}$
for all $x$ in the domain $D(M)=\left\{x\in X\,\middle|\,M(x)\neq\emptyset\right\}$. The range of $M$ is defined by $R(M)=\cup_{x\in D(M)}M(x)$. To avoid confusion, we denote the zeros of a nonlinear possibly set-valued operator $M:X\rightrightarrows X$ by
$\operatorname{zer}M\coloneqq\left\{x\in X\,\middle|\,0\in M(x)\right\}=M^{-1}(\{0\}).$
A set-valued operator $M:X\rightrightarrows X$ is called monotone if
$\langle y_{1}-y_{2},x_{1}-x_{2}\rangle_{X}\geq 0$ (2.1)
for all $x_{1},x_{2}\in D(M)$ and $y_{1}\in M(x_{1}),y_{2}\in M(x_{2})$. The operator $M$ is maximally monotone (m-monotone), if $M$ is monotone and has no proper monotone extension (equivalently, $R(\lambda I+M)=X$ holds for some (and hence for all) $\lambda>0$, [1]). A prominent example of maximal monotone operators is given by subgradients of proper functions $F:X\to\mathbb{R}\cup\{\infty\}$. The subgradient of $F$ at $x_{0}\in X$ is the set-valued operator $\partial F:X\rightrightarrows X^{*}$ with
$\partial F(x_{0})=\left\{x^{\prime}\in X^{*}\,\middle|\,\langle x^{\prime},x-x_{0}\rangle_{X^{*},X}\leq F(x)-F(x_{0})\,\forall\,x\in X\right\},$
see Appendix 8.2.
Definition 2.1.
Let $(X,\langle\cdot,\cdot\rangle_{X})$ and $(U,\langle\cdot,\cdot\rangle_{U})$ be real Hilbert spaces, let $M$ be maximally monotone and $B\in L(U,X)$. Then, a maximally monotone pH system is defined by the differential expression
$\begin{split}\tfrac{\mathrm{d}}{\mathrm{d}t}x(t)+M(x(t))&\ni Bu(t),\\ y(t)&=B^{*}x(t),\end{split}$ (mpHS)
with $x(0)=x_{0}\in\overline{D(M)}$ and where $B^{*}\in L(X,U)$ denotes the Hilbert space adjoint of the bounded operator $B$.
For ease of notation, we will occasionally write $(M,B)$ when referring to the system (mpHS).
We provide a brief overview w.r.t well-posedness of maximally monotone pH systems. To this end, we first consider the homogeneous Cauchy problem, that is,
$\tfrac{\mathrm{d}}{\mathrm{d}t}x(t)+M(x(t))\ni 0,\quad x(0)=x_{0}.$ (2.2)
For this autonomous problem, for every $x_{0}\in\overline{D(M)}$ a unique mild solution $x\in C([0,t_{f}],X)$ in the sense of [1, Def. 4.3] may be defined by the exponential formula
$T(t)x_{0}=\lim_{n\rightarrow\infty}\left(I+\tfrac{t}{n}M\right)^{-n}x_{0}.$ (2.3)
By the maximal monotonicity of $M$, it follows directly from [1, Definition 4.4], that the the exponential formula (2.3) defines a nonexpansive semigroup $(T(t))_{t\geq 0}:\overline{D(M)}\to\overline{D(M)}$ of nonlinear operators, i.e.,
$\|T(t)x-T(t)y\|_{X}\leq\|x-y\|_{X}$
for all $t\geq 0$, $x,y\in\overline{D(M)}$.
For the inhomogeneous case, let $t_{f}>0$ and $f\in L^{1}([0,t_{f}],X)$. Then, for all $x_{0}\in D(M)$ there exists a unique strong solution $x\in W^{1,\infty}([0,t_{f}],X)$ of the inhomogeneous Cauchy problem
$\tfrac{\mathrm{d}}{\mathrm{d}t}x(t)+M(x(t))\ni f(t),\quad x(0)=x_{0},$ (2.4)
by [1, Theorem 4.4]. Moreover, the unique strong solution fulfills (2.4) for almost every $t\in[0,t_{f}]$. As $B\in L(U,X)$, this implies that for all $u\in L^{1}([0,t_{f}],U)$ and $x_{0}\in D(M)$ there exists a unique strong solution $x\in W^{1,\infty}([0,t_{f}],X)$ of (mpHS).
2.1. Shifted passivity and stability
Define the principal section $M_{0}:D(M)\to X$ of $M$ via
$M_{0}(x)=\underset{y\in M(x)}{\operatorname{arg\,min}}\|y\|_{X}^{2},\qquad x\in D(M).$ (2.5)
Note that, in contrast to $M$, $M_{0}$ is a single-valued operator due to maximality of $M$ [25]. It is well-known, see again [25], that the right (strong) derivative at $t$
$\frac{\mathrm{d}^{+}}{\mathrm{d}t}x(t)\;:=\;\lim_{h\downarrow 0}\frac{1}{h}x(t+h)-x(t),$
of the unique strong solution $x\in W^{1,\infty}([0,t_{f}],X)$ of (mpHS), exists everywhere in $\mathbb{R}_{+}$ and satisfies
$\frac{\mathrm{d}^{+}}{\mathrm{d}t}x(t)=-M_{0}(x(t))+Bu(t)$ (2.6)
almost everywhere in $0,t_{f}]$, see [3]. Since the maximal monotonicity of $M$ implies by definition (2.5) that also $M_{0}$ is maximally monotone, the system (2.6) defines a maximally monotone pH system in the sense of [10].
As in the case of single-valued operators, see [10, Prop. 2.3], the system (mpHS) may be shown to be passive with respect to steady state pairs. Here, we call $(\overline{x},\overline{u})\in D(M)\times U$ a steady state pair of (mpHS) if
$M(\overline{x})\ni B\overline{u}.$ (2.7)
Proposition 2.2.
Let $(M,B)$ be a maximally monotone pH system and $(\overline{x},\overline{u})\in D(M)\times U$ a steady state pair of $(M,B)$. Then, for all $u\in L^{1}([0,t_{f}],U)$ and for almost every $t\in[0,t_{f}]$ we have the passivity inequality
$\tfrac{1}{2}\tfrac{\mathrm{d}^{+}}{\mathrm{d}t}\|x(t)-\overline{x}\|_{X}^{2}$ $=-\langle x(t)-\overline{x},M_{0}(x(t))-M_{0}(\overline{x})\rangle_{X}+\langle y(t)-\overline{y},u(t)-\overline{u}\rangle_{U}$
$\leq\langle y(t)-\overline{y},u(t)-\overline{u}\rangle_{U},$
where $\overline{y}=B^{*}\overline{x}$ is the steady-state output.
Proof.
As constant trajectory $\overline{x}$ solves (mpHS),
$0=\frac{\mathrm{d}^{+}}{\mathrm{d}t}\overline{x}=-M_{0}(\overline{x})+B\overline{u}$
and hence, we have
$\tfrac{\mathrm{d}}{\mathrm{d}t}(x(t)-\bar{x})=-M_{0}(x(t))+M_{0}(\bar{x})+B(u(t)-\bar{u})$
for any solution of (mpHS). Since $M_{0}$ is also maximally monotone, this directly implies the desired passivity inequality. ∎
2.2. Structure-preserving interconnection
We now prove that the presented class of nonlinear pH systems subject to set-valued operators is closed under power-preserving interconnections.
Proposition 2.3.
Let $(M_{1},B_{1})$ and $(M_{2},B_{2})$ be two (maximally) monotone pH systems in the Hilbert spaces $X_{1}$ and $X_{2}$, with inputs and outputs $u_{i},y_{i}\in U_{i}$, $i=1,2$. Assume that $U_{i}$ is decomposed as $U_{i}=U_{i}^{1}\times U_{i}^{2}$ such that $u_{i}$ and $y_{i}$ admit a splitting $u_{i}=(u_{i}^{1},u_{i}^{2})$ and $y=(y_{i}^{1},y_{i}^{2})$, respectively. Correspondingly, the input maps are represented as $B_{i}=\left[\begin{smallmatrix}B_{i}^{1}&B_{i}^{2}\end{smallmatrix}\right]$, $i=1,2$. For $\mathcal{E}\in L(U_{2}^{1},U_{1}^{1})$ define
$\begin{bmatrix}0&\mathcal{E}\\ -\mathcal{E}^{*}&0\end{bmatrix}\begin{bmatrix}y_{1}^{1}\\ y_{2}^{1}\end{bmatrix}=\begin{bmatrix}u_{1}^{1}\\ u_{2}^{1}\end{bmatrix}$. (2.8)
Then, the interconnected system
[…]
where
[…]
defines a (maximally) monotone pH system.
Proof.
Note that $M$ may be decomposed into $M=N+K$ where
$N=\begin{bmatrix}M_{1}\\ M_{2}\end{bmatrix}\quad\mathrm{and}\quad K=\begin{bmatrix}0&-B_{1}^{1}\mathcal{E}(B_{2}^{1})^{*}\\ B_{2}^{1}\mathcal{E}^{*}(B_{1}^{1})^{*}&0\end{bmatrix}$.
Monotonicity of $M$ follows directly since $N$ is monotone and $K$ is skew-adjoint. Further, maximality of $M_{1}$ and $M_{2}$ imply maximality of $N$. Then, the sum rule for maximally monotone operators [2, Cor. 25.5] yields that $M$ is maximally monotone because $D(K)=X_{1}\times X_{2}$ due to boundedness of $B_{i}$, $i=1,2$ and $\mathcal{E}$. ∎
The interconnection law (2.8) is not the only one that preserves the monotonicity structure. In what follows, we consider a second important class of suitable couplings induced by nonlinear and nonexpansive operators which will be particularly useful when considering control-constrained stabilization later.
Let $D\subset X$ be a nonempty subset. A (possibly nonlinear) operator $\mathcal{E}:D\to X$ is called firmly nonexpansive if
$\langle\mathcal{E}(x)-\mathcal{E}(y),x-y\rangle\geq\|\mathcal{E}(x)-\mathcal{E}(y)\|^{2}$ (2.10)
for all $x,y\in D$, see [2, Prop. 4.4].
Proposition 2.4.
Let $(M_{1},B_{1})$ and $(M_{2},B_{2})$ be two (maximally) monotone pH systems in the Hilbert spaces $X_{1}$ and $X_{2}$, with inputs and outputs $u_{i},y_{i}\in U_{i}$, $i=1,2$. Let $H$ be another Hilbert space, $\mathcal{E}:H\to H$ be firmly nonexpansive and $\mathcal{F}_{i}\in L(U_{i},H)$, $i=1,2$. Define the coupling
$\begin{bmatrix}u_{1}\\ u_{2}\end{bmatrix}=\begin{bmatrix}-c&0\\ 0&c\end{bmatrix}\begin{bmatrix}\mathcal{F}_{1}^{*}\mathcal{E}(\mathcal{F}_{1}y_{1})-\mathcal{F}_{1}^{*}\mathcal{E}(\mathcal{F}_{2}y_{2})\\ \mathcal{F}_{2}^{*}\mathcal{E}(\mathcal{F}_{1}y_{1})-\mathcal{F}_{2}^{*}\mathcal{E}(\mathcal{F}_{2}y_{2})\end{bmatrix}$. (2.11)
for a positive constant $c>0$. Then, the autonomous interconnected system
$\tfrac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}+M\begin{bmatrix}x_{1}(t)\\ x_{2}(t)\end{bmatrix}$ $\ni 0$ (2.12)
defines a (maximally) monotone pH system, where
[…]
with
[…]
Proof.
We decompose $M=N+K$ where $N=\left[\begin{smallmatrix}M_{1}\\ M_{2}\end{smallmatrix}\right]$. Since $N$ is maximally monotone it is, again due to the sum rule for maximally monotone operators, enough to prove that $K$ is monotone and $D(K)=X_{1}\times X_{2}$. To this end, consider for $x_{i}\in D(M_{i})$, $i=1,2$:
$\left\langle K\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix},\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}\right\rangle_{X_{1}\times X_{2}}$ $\quad=c\left\langle\begin{bmatrix}\mathcal{E}(\mathcal{F}_{1}B_{1}^{*}x_{1})-\mathcal{E}(\mathcal{F}_{2}B_{2}^{*}x_{2})\\ \mathcal{E}(\mathcal{F}_{1}B_{1}^{*}x_{1})-\mathcal{E}(\mathcal{F}_{2}B_{2}^{*}x_{2})\end{bmatrix},\left[\begin{smallmatrix}\mathcal{F}_{1}B_{1}^{*}x_{1}\\ -\mathcal{F}_{2}B_{2}^{*}x_{2}\end{smallmatrix}\right]\right\rangle_{H\times H}$
$\quad=c\langle\mathcal{E}(\mathcal{F}_{1}B_{1}^{*}x_{1})-\mathcal{E}(\mathcal{F}_{2}B_{2}^{*}x_{2}),\mathcal{F}_{1}B_{1}^{*}x_{1}-\mathcal{F}_{2}B_{2}^{*}x_{2}\rangle_{H}$
$\quad\geq c\|\mathcal{E}(\mathcal{F}_{1}B_{1}^{*}x_{1})-\mathcal{E}(\mathcal{F}_{2}B_{2}^{*}x_{2})\|_{H}^{2}\geq 0,$
where the first inequality follows by firm nonexpansiveness of $\mathcal{E}$, [2, Prop. 4.2]. Hence, $K$ is monotone, and its maximality is a consequence of $\mathcal{E}$ and $\mathcal{F}_{i}$ being everywhere defined. ∎
3. Monotonicity in optimal control
Before developing the control-by-interconnection scheme, we examine the monotone structures inherent in the optimality conditions of possibly inequality-constrained optimal control problems. To this end, define $X\coloneqq H^{1}([0,t_{f}],\mathbb{R}^{n})$, $P\coloneqq L^{2}([0,t_{f}],\mathbb{R}^{n})$ and $U\coloneqq L^{2}([0,t_{f}],\mathbb{R}^{m})$. We denote the continuous embedding from $X$ to $P$ by $\iota:X\hookrightarrow P$.
We consider the optimal control problem on the finite horizon $t_{f}>0$
$\begin{aligned} \min_{(x,u)\in P\times U}\quad&J(x,u)\\ \text{s.t.}\quad&\dot{x}(\tau)=Ax(\tau)+Bu(\tau)+f(\tau),\quad x(0)=x_{0},\\ &\underline{u}\leq u(t)\leq\overline{u}\quad\text{a.e. on }[0,t_{f}].\end{aligned}$ (3.1)
where $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$ is injective, $f\in P$, $x_{0}\in\mathbb{R}^{n}$, $\underline{u},\overline{u}\in\mathbb{R}^{m}$ with $\underline{u}<\overline{u}$ componentwise and the cost functional $J:P\times U\to\mathbb{R}$ is defined by
$J(x,u):=\int_{0}^{t_{f}}\ell_{x}(x(\tau))+\ell_{u}(u(\tau))\,\mathrm{d}\tau$ (3.2)
where $\ell_{x}:\mathbb{R}^{n}\rightarrow[0,\infty)$ is continuous and convex and $\ell_{u}:\mathbb{R}^{m}\rightarrow[0,\infty)$ continuous and strictly convex.
We further assume throughout that the $L^{2}$-gradient $\nabla J(x,u)$ of the cost functional $J$ exists. To derive optimality conditions of (3.1) we introduce the following operators and sets. Denote the densely-defined unbounded operator
$\mathcal{C}$ $:P\times U\supset D(\mathcal{C})\to P\times\mathbb{R}^{n},$
$D(\mathcal{C})$ $\coloneqq X\times U,\qquad\mathcal{C}\begin{bmatrix}x\\ u\end{bmatrix}\coloneqq\begin{bmatrix}\tfrac{\mathrm{d}}{\mathrm{d}\tau}x-Ax-Bu\\ x(0)\end{bmatrix}$ (3.3)
where $(x,u)\mapsto Ax+Bu$ is viewed as a linear and bounded multiplication operator on $P\times U$. We recall from [10, Lem. 2.3] that its unbounded Hilbert space adjoint $\mathcal{C}^{*}$ is given by
$\mathcal{C}^{*}:P\times\mathbb{R}^{n}\supset D(\mathcal{C}^{*})\to P\times U$
where
$D(\mathcal{C}^{*})=\{(\lambda,\lambda_{0})\in X\times\mathbb{R}^{n}\,|\,\lambda(0)=\lambda_{0}\wedge\lambda(t_{f})=0\}$
and
$\mathcal{C}^{*}\begin{bmatrix}\lambda\\ \lambda_{0}\end{bmatrix}=\begin{bmatrix}-\frac{\mathrm{d}}{\mathrm{d}\tau}\lambda-A^{\top}\lambda\\ -B^{\top}\lambda\end{bmatrix}$.
In addition, let
$F\coloneqq\left\{u\in U\,\middle|\,\underline{u}_{i}\leq u_{i}(t)\leq\overline{u}_{i}\ \text{a.e. on}\ [0,t_{f}]\right\}$ (3.4)
such that the optimal control problem (3.1) may be written compactly as
$\min_{(x,u)\in P\times U}J(x,u)\quad\mathrm{s.t.}\quad\mathcal{C}\begin{bmatrix}x\\ u\end{bmatrix}=\begin{bmatrix}f\\ x_{0}\end{bmatrix},\ u\in F.$
The following result shows that the optimality system is governed by the set-valued and maximally monotone mapping
$M_{\mathrm{opt}}:\begin{bmatrix}\left[\begin{smallmatrix}P\\ U\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}P\\ \mathbb{R}^{n}\end{smallmatrix}\right]\\ U\end{bmatrix}\supset\begin{bmatrix}D(\mathcal{C})\\ D(\mathcal{C}^{*})\\ U\end{bmatrix}\rightrightarrows\begin{bmatrix}\left[\begin{smallmatrix}P\\ U\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}P\\ \mathbb{R}^{n}\end{smallmatrix}\right]\\ U\end{bmatrix}\eqqcolon\mathcal{W}$ (3.5)
with
$\underbrace{\begin{bmatrix}\left[\begin{smallmatrix}x\\ u\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}\lambda\\ \lambda_{0}\end{smallmatrix}\right]\\ \mu\end{bmatrix}}_{\eqqcolon w}\mapsto\underbrace{\begin{bmatrix}\nabla J(x,u)\\ 0\\ \partial\sigma_{F}(\mu)\end{bmatrix}}_{\eqqcolon\mathcal{M}_{1}(w)}+\underbrace{\begin{bmatrix}0&\mathcal{C}^{*}&\left[\begin{smallmatrix}0\\ I\end{smallmatrix}\right]\\ -\mathcal{C}&0&0\\ \left[\begin{smallmatrix}0\hskip 1.42262pt,\hskip 2.84526pt-I\end{smallmatrix}\right]&0&0\end{bmatrix}}_{\eqqcolon\mathcal{M}_{2}}\!\!\begin{bmatrix}\left[\begin{smallmatrix}x\\ u\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}\lambda\\ \lambda_{0}\end{smallmatrix}\right]\\ \mu\end{bmatrix}$,
where $\sigma_{C}(x)=\sup_{c\in C}\langle c,x\rangle_{X}$ denotes the support function, see Appendix 8.2.
Theorem 3.1.
The optimal control problem (3.1) has a unique solution $(x^{\star},u^{\star})\in D(\mathcal{C})$. Moreover, there exist unique $(\lambda,\lambda_{0})\in D(\mathcal{C}^{*}),\mu\in U$ such that
[…]
Moreover, the set-valued operator $M_{\mathrm{opt}}:\mathcal{W}\rightrightarrows\mathcal{W}$ given by (3.5) is maximally monotone. The resolvents $(\lambda I+M_{\mathrm{opt}})^{-1}$ are firmly nonexpansive for all $\lambda>0$.
Proof.
By using the same argumentation as in [10, Prop. 3.4], the existence and uniqueness of solutions to (3.1) is immediate from [9, Prop. 2.1]. Define the operator
$C:X\times U\to P\times\mathbb{R}^{n},\quad C\begin{bmatrix}x\\ u\end{bmatrix}\coloneqq\begin{bmatrix}\tfrac{\mathrm{d}}{\mathrm{d}\tau}x-Ax-Bu\\ x(0)\end{bmatrix}$
which is the operator $\mathcal{C}$ considered on the Sobolev space $X$, and therefore bounded, and let $z_{0}\in X\times U$ be such that $Cz_{0}=\left[\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right]$. To penalize the infeasibility of $(x,u)$ with infinite costs we introduce $f:P\times U\to\mathbb{R}\cup\{\infty\}$ via
$f(x,u)=J(\iota x,u)+i_{\ker C+\{z_{0}\}}(x,u)+i_{X\times F}(x,u).$
where $i_{K}:Y\to\mathbb{R}\cup\{+\infty\}$ denotes the indicator function
$i_{K}(y)\coloneqq\begin{cases}0,&\text{if }y\in K,\\ +\infty,&\text{otherwise}.\end{cases}$
of some subset $K$ of the underlying Hilbert space $Y$. This way, the problem (3.1) is equivalent to
$\min_{(x,u)\in P\times U}f(x,u)$
and we obtain
$(x^{\star},u^{\star})=\underset{(x,u)\in X\times U}{\operatorname{arg\,min}}f=(\partial f)^{-1}(\{0\})$
from [2, Thm. 16.3]. In other words, for the unique minimizer $(x^{\star},u^{\star})\in X\times U$ of $f$ it holds that
$\partial f(x^{\star},u^{\star})\ni 0.$
We denote the domain of a function $f:Y\to\overline{R}$ by
$\operatorname{dom}(f)\coloneqq\left\{y\in Y\,\middle|\,f(y)<+\infty\right\}.$
Since $\operatorname{dom}(J)=X\times U$, [2, Cor. 16.48] yields that
$\partial f(x^{\star},u^{\star})=\partial J(x^{\star},u^{\star})+\partial(i_{\ker C+\{z_{0}\}}+i_{X\times F})(x^{\star},u^{\star}).$
Furthermore, we have that $\operatorname{dom}(i_{\ker C+\{z_{0}\}})=\ker C+\{z_{0}\}$ and $\operatorname{int}\operatorname{dom}(i_{X\times F})=X\times\operatorname{int}F.$ Clearly, by assumption $0\in\operatorname{int}F$. By solvability of the state equation for any $u$, there is an associated state $x$ such that $\mathcal{C}\left(\begin{smallmatrix}x\\ 0\end{smallmatrix}\right)=\left(\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right)$. Thus,
$\left(\begin{smallmatrix}x\\ 0\end{smallmatrix}\right)\in\operatorname{dom}(i_{\ker C+\{z_{0}\}})\cap\operatorname{int}\operatorname{dom}(i_{X\times F})$
such that the right-hand side is non-empty. Hence, again using [2, Cor. 16.48] we get
$\partial f(x^{\star},u^{\star})=\partial J(x^{\star},u^{\star})+\partial i_{\ker C+\{z_{0}\}}(x^{\star},u^{\star})+\partial i_{X\times F}(x^{\star},u^{\star}).$
Moreover, the $L^{2}$-gradient $\nabla J(x,u)$ of the cost functional $J$ exists by assumption and with that we have $\partial J(x^{\star},u^{\star})=\{\nabla J(x^{\star},u^{\star})\}$, [2, Prop. 17.31]. Since $(x^{\star},u^{\star})$ is a minimizer of $f$, it is feasible and hence, $(x^{\star},u^{\star})\in\ker C+\{z_{0}\}$. With that, we use [2, Ex. 16.13] and Proposition 8.3 to obtain
$\partial i_{\ker C+\{z_{0}\}}\left(\begin{smallmatrix}x^{\star}\\ u^{\star}\end{smallmatrix}\right)=N_{\ker C}\left(\begin{smallmatrix}x^{\star}\\ u^{\star}\end{smallmatrix}\right)=(\ker C)^{\perp}=R(C^{*}),$
where the third equality is a consequence of the closed range theorem and [10, Lem. 3.2]. Finally,
$\partial i_{X\times F}(x^{\star},u^{\star})=\left[\begin{smallmatrix}\{0_{X^{*}}\}\\ \partial i_{F}(u^{\star})\end{smallmatrix}\right].$
Hence, there exist $(\lambda,\lambda_{0})\in P\times\mathbb{R}^{n}$ and $\mu\in\partial i_{F}(u^{\star})$ such that
$C^{*}\left(\begin{smallmatrix}\lambda\\ \lambda_{0}\end{smallmatrix}\right)=-\nabla J\left(\begin{smallmatrix}x^{\star}\\ u^{\star}\end{smallmatrix}\right)-\left[\begin{smallmatrix}0\\ \mu\end{smallmatrix}\right]\in P\times U$
and with that $(\lambda,\lambda_{0})\in D(\mathcal{C}^{*})$. Consequently, there exist $(\lambda,\lambda_{0})\in D(\mathcal{C}^{*})$ and $\mu\in\partial i_{F}(u^{\star})$ such that
$\nabla J\left(\begin{smallmatrix}x^{\star}\\ u^{\star}\end{smallmatrix}\right)+\mathcal{C}^{*}\left(\begin{smallmatrix}\lambda\\ \lambda_{0}\end{smallmatrix}\right)+\left[\begin{smallmatrix}0\\ \mu\end{smallmatrix}\right]=\left[\begin{smallmatrix}0\\ 0\end{smallmatrix}\right]\in P\times U$ (adjoint)
holds. By feasibility of $(x^{\star},u^{\star})$ we obtain
$0=\mathcal{C}(z_{0}-\left(\begin{smallmatrix}x^{\star}\\ u^{\star}\end{smallmatrix}\right))=\left[\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right]-\mathcal{C}\left(\begin{smallmatrix}x^{\star}\\ u^{\star}\end{smallmatrix}\right).$ (state)
Following [9, Cor. 5.2], we have the duality
$\mu\in\partial i_{F}(u^{\star})\quad\Longleftrightarrow\quad u^{\star}\in\partial i_{F}^{*}(\mu),$ (3.7)
where $i_{F}^{*}=\sigma_{F}$ denotes the Fenchel conjugate of $i_{F}$, see Proposition 8.2. As a consequence, (3.7) yields the feasibility condition
$\partial\sigma_{F}(\mu)-\{u^{\star}\}\ni 0.$ (feasibility)
In summary, there exist $(\lambda,\lambda_{0})\in D(\mathcal{C}^{*}),\mu\in U$ such that
[…]
It remains to show that $M_{\mathrm{opt}}$ is an injection. To this end, let $w_{i}=(x_{i},u_{i},\lambda_{i},\lambda_{0i},\mu_{i})\in D(M_{\mathrm{opt}}),i=1,2$, such that $M_{\mathrm{opt}}w_{1}=M_{\mathrm{opt}}w_{2}$. Then,
$0$ $=\langle M_{\mathrm{opt}}(w_{1})-M_{\mathrm{opt}}(w_{2}),w_{1}-w_{2}\rangle_{\mathcal{W}}$
$=\langle\mathcal{M}_{1}(w_{1})-\mathcal{M}_{2}(w_{2}),w_{1}-w_{2}\rangle_{\mathcal{W}}$
$=\langle\nabla J(x_{1},u_{1})-\nabla J(x_{2},u_{2}),(x_{1},u_{1})-(x_{2},u_{2})\rangle_{P\times U}$
$\hphantom{=}+\langle\partial\sigma_{F}(\mu_{1})-\partial\sigma_{F}(\mu_{2}),\mu_{1}-\mu_{2}\rangle_{U}$
$\geq\alpha\|u_{1}-u_{2}\|^{2}_{U}$
for some $\alpha>0$. Hence, $u_{1}=u_{2}$ and which implies $x_{1}=x_{2}$ by means of the variation of parameters formula (8.1). Consequently,
$\mathcal{C}^{*}\left(\begin{bmatrix}\lambda_{1}\\ \lambda_{01}\end{bmatrix}-\begin{bmatrix}\lambda_{2}\\ \lambda_{02}\end{bmatrix}\right)+\begin{bmatrix}0\\ \mu_{1}-\mu_{2}\end{bmatrix}=0$
shows that $\begin{bmatrix}\lambda_{1}\\ \lambda_{01}\end{bmatrix}=\begin{bmatrix}\lambda_{2}\\ \lambda_{02}\end{bmatrix}$ since in particular
$-\frac{\mathrm{d}}{\mathrm{d}\tau}-A^{\top}:P\times\mathbb{R}^{n}\supset D(\mathcal{C}^{*})\to P$
is injective, [10, Lem. 3.2]. Hence, $\mu_{1}=\mu_{2}$ which proves injectivity of $M_{\mathrm{opt}}$.
We now proceed with the maximal monotonicity of $M_{\mathrm{opt}}$. Using the notation that is introduced below (3.5), we have $M_{\mathrm{opt}}=\mathcal{M}_{1}+\mathcal{M}_{2}$. Here $\mathcal{M}_{1}$ is the subgradient of the proper lower semicontinuous convex function $m_{1}(w)=J(x,u)+\sigma_{F}(\mu)$ and thus, maximally monotone. The operator $\mathcal{M}_{2}$ is skew-adjoint and with that also maximally monotone. Hence, $M_{\mathrm{opt}}$, as the sum of two monotone operators, is monotone. It remains to show that $M_{\mathrm{opt}}$ is maximal. Following [2, Cor. 25.5], it suffices to prove $D(\mathcal{M}_{1})=\mathcal{W}$. Observe that
$D(\mathcal{M}_{1})$ $=\left\{w\in\mathcal{W}\,\middle|\,\mathcal{M}_{1}(w)\neq\emptyset\right\}=\left[\begin{smallmatrix}P\\ U\end{smallmatrix}\right]\times\left[\begin{smallmatrix}P\\ \mathbb{R}^{n}\end{smallmatrix}\right]\times D(\partial\sigma_{F})$
which proves that only the full domain of $\partial\sigma_{F}$ must be verified. Define the set-valued operator $T:U\rightrightarrows U$ via
$T(\mu)\coloneqq\left\{f\in F\,\middle|\,\langle f,\mu\rangle_{U}=\sigma_{F}(\mu)\right\},\quad\mu\in U.$
We show that $T\subset\partial\sigma_{F}$ and $D(T)=U$. To this end, let $\mu\in U$ and $f\in T(\mu)$ which yields $f\in U$ and
$\langle f,z-\mu\rangle_{U}\leq\sup_{f\in F}\langle f,z\rangle_{U}-\langle f,\mu\rangle_{U}=\sigma_{F}(z)-\sigma_{F}(\mu)$
for all $z\in U$. Consequently, $f\in U$ is a subdifferential of $\sigma_{F}$ at $\mu$ and we conclude $D(T)\subset D(\partial\sigma_{F})$. Thus, we remain to prove $D(T)=U$. We now fix an arbitrary $\mu\in U$. Then, there exists a maximizing sequence $(f_{n})_{n\in\mathbb{N}}\subset F$ that fulfills
$\lim_{n\to\infty}\langle\mu,f_{n}\rangle_{U}=\sup_{f\in F}\langle\mu,f\rangle_{U}=\sigma_{F}(\mu)<\infty,$
where the last inequality follows as $F$ is bounded. Since $F$ is a bounded subset of a the Hilbert space $U$, $F$ is relatively weakly sequentially compact, [23, Thm. 2.10]. Therefore, $(f_{n})_{n\in\mathbb{N}}$ admits a weakly convergent subsequence $(f_{n_{k}})_{k\in\mathbb{N}}\subset F$ with
$f_{n_{k}}\rightharpoonup f^{\star}\in U$
as $k\to\infty$. But $F$ is closed and convex and with that, $F$ is weakly sequentially closed, [23, Thm. 2.11], which proves that $f^{\star}\in F$. Hence,
$\langle\mu,f^{\star}\rangle_{U}=\lim_{k\to\infty}\langle\mu,f_{n_{k}}\rangle_{U}=\lim_{n\to\infty}\langle\mu,f_{n}\rangle_{U}=\sigma_{F}(\mu)$
which gives $U=D(T)\subset D(\partial\sigma_{F})\subset U$. Eventually, $D(\partial\sigma_{F})=U$. Thus, $M_{\mathrm{opt}}$ is maximally monotone which yields that the resolvents $J_{\lambda}\coloneqq(\lambda I+M_{\mathrm{opt}})^{-1}$ are single-valued everywhere defined firmly nonexpansive mappings in $\mathcal{W}$ [2, Prop. 23.8].∎
4. Stability of open-loop primal-dual gradient dynamics
In this part, we formulate a primal-dual gradient method to iteratively solve the optimal control problem (3.1) or equivalently the optimality conditions of Theorem 3.1. Abstractly, performing gradient ascent in the dual variable and gradient descent in the primal variable, the method is given by the monotone pHs
$\tfrac{\mathrm{d}}{\mathrm{d}t}w(t)+M_{\mathrm{opt}}(w(t))\ni\mathcal{B}u_{\mathrm{opt}}(t),$ (4.1)
with input operator defined by
$\mathcal{B}\coloneqq\begin{bmatrix}0\\ -I_{2n}\\ 0\end{bmatrix}$.
This input operator will encode the data provided to the optimal control problem by the plant given by the pair $(x_{0},f)$. In particular, it allows to set the initial value in view of coupling with the plant, see Figures 1 and 2.
Due to Theorem 3.1, (4.1) defines a maximally monotone pH system with unique controlled steady state given by
$(w^{\star},u_{\mathrm{opt}}^{\star})=((x^{\star},u^{\star},\lambda,\lambda_{0},\mu),(f,x_{0})),$ (4.2)
and hence provides a one-to-one correspondence of (4.1) and optimal solutions. The following corollary is a direct consequence of Proposition 2.2 and provides a shifted-passivity-type estimate for the current iterate.
Corollary 4.1.
For all $u_{\mathrm{opt}}\in L^{1}([0,t_{f}],P\times\mathbb{R}^{n})$ and for almost every $t\in[0,t_{f}]$ it holds that
$\tfrac{1}{2}\tfrac{\mathrm{d}^{+}}{\mathrm{d}t}\|w(t)-w^{\star}\|^{2}\leq\langle y_{\mathrm{opt}}(t)-\mathcal{B}^{*}w^{\star},u_{\mathrm{opt}}(t)-\left(\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right)\rangle_{U}$
where $w\in W^{1,\infty}([0,t_{f}],\mathcal{W})$ denotes the unique strong solution to (4.1). In particular, if the input fulfills $u_{\mathrm{opt}}(t)\equiv\left(\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right)$, we have that the mapping $t\mapsto\|w(t)-w^{\star}\|^{2}$ is monotonically nonincreasing.
Before proving a stability result in the next subsection, we briefly mention that finite-dimensional saddle-point dynamics occuring in optimal control have been analyzed in the works [6, 7]. Here, as we consider the continuous-time optimal control problem, the dynamics are Hilbert space valued.
4.1. Stability of unconstrained optimizer dynamics
In infinite-dimensional spaces it can be challenging to determine the asymptotic behavior of nonlinear systems such as (2.2). A well-established tool are LaSalle invariance principles which we will leverage to prove stability of the continuous-time primal-dual gradient method given by (4.1).
LaSalle-type invariance principle. Consider the nonlinear homogeneous Cauchy problem (2.2) with densely-defined $M$, i.e., $\overline{D(M)}=X$. Let $(T_{M}(t))_{t\geq 0}:X\to X$ be the family of nonlinear contractions defined by (2.3). Let $V:X\to\mathbb{R}_{+}$ be a weakly lower semi-continuous functional. For each $x_{0}\in X$ we define the mapping $V_{M}(\cdot,x_{0}):\mathbb{R}_{+}\to\mathbb{R}_{+}$ via
$t\mapsto V(T_{M}(t)x_{0})\eqqcolon V_{M}(t,x_{0}).$
For fixed $x_{0}\in X$ and $s\in\mathbb{R}_{+}$, define the upper Dini derivative of $V_{M}(\cdot,x_{0})$ at $s$ by
$\dot{V}_{M}(s,x_{0})\coloneqq\limsup_{t\downarrow 0}\frac{1}{t}(V_{M}(s+t,x_{0})-V_{M}(s,x_{0})).$
We say that $\dot{V}_{M}(s,x_{0})$ exists if the above limit is finite. We call $V:X\to\mathbb{R}_{+}$ a Lyapunov functional for (2.2), if $\dot{V}_{M}(0,x_{0})$ exists for all $x_{0}\in X$
$\dot{V}_{M}(x_{0})\coloneqq\dot{V}_{M}(0,x_{0})\leq 0.$
The proof of the following theorem can be found in [21, Thm. 2.3].
Theorem 4.2 (Weak invariance principle).
Let $Z\subset X$ be weakly sequentially compact with respect to the topology in $X$. Moreover, let $Z$ be positively invariant with respect to (2.2). Let $V:X\to\mathbb{R}$ be weakly lower semi-continuous Lyapunov functional for (2.2). Define
$S_{0}\coloneqq\left\{z_{0}\in Z\,\middle|\,\dot{V}_{M}(z_{0})=0\right\}$ (4.3)
and assume that the largest invariant set in $S_{0}$ is a singleton $\{s_{0}\}$. Then, for all $z_{0}\in Z$ we have
$T_{M}(t)x_{0}\rightharpoonup s_{0},\quad\text{ as $t\to\infty$.}$
The governing operators of primal-dual gradient dynamics given in (4.1) admit a very specific structure. Precisely, the saddle-point structure causes a nonlinear everywhere defined monotone perturbation of an unbounded skew-adjoint operator. As we will make use of Theorem 4.2 our aim is to give a concrete representation of the level sets $S_{0}$ as in (4.3). We define the functional $V:X\to\mathbb{R}$ via
[…]
Theorem 4.3.
Let $\mathfrak{A}:X\supset D(\mathfrak{A})\to X$ be a densely defined and skew-adjoint operator and $\mathfrak{B}:X\to X$ be m-monotone and continuous. Then, the level sets of $\dot{V}_{\mathfrak{A}+\mathfrak{B}}$ are given by
$\dot{V}_{\mathfrak{A}+\mathfrak{B}}^{-1}(\{c\})=\left\{x\in X\,\middle|\,-\langle\mathfrak{B}(x),x\rangle_{X}=c\right\}.$
Proof.
The proof is given in Appendix 8.1. ∎
The subsequent theorem generalizes the last result to Lyapunov functionals that are shifted with respect to the steady states of the underlying dynamics.
Theorem 4.4.
Let $\mathfrak{A}:X\supset D(\mathfrak{A})\to X$ be a densely defined and skew-adjoint operator and let $\mathfrak{B}:X\to X$ be m-monotone and continuous. Let $(U(t))_{t\geq 0}$ denote the strongly continuous semigroup of nonexpansive mappings generated by $(\mathfrak{A}+\mathfrak{B})$. Assume there exists a unique $s_{0}\in\operatorname{zer}(\mathfrak{A}+\mathfrak{B})$, i.e.,
$\mathfrak{A}s_{0}+\mathfrak{B}(s_{0})=0.$ (4.5)
Define the functional $V_{s_{0}}:X\to\mathbb{R}$ via
$V_{s_{0}}(x)\coloneqq\tfrac{1}{2}\|x-s_{0}\|_{X}^{2}.$ (4.6)
Then the upper Dini derivative $\dot{V}_{s_{0},\mathfrak{A}+\mathfrak{B}}:X\to\mathbb{R}$ given by
$\dot{V}_{s_{0},\mathfrak{A}+\mathfrak{B}}(x)\coloneqq\limsup_{t\downarrow 0}\frac{1}{2t}\Big(\|U(t)(x)-s_{0}\|_{X}^{2}-\|x-s_{0}\|_{X}^{2}\Big)$ (4.7)
has level sets given by
$\dot{V}_{s_{0},\mathfrak{A}+\mathfrak{B}}^{-1}(\{c\})$ (4.8)
$=\Big\{x\in X\,\Big|\,-\big\langle\mathfrak{B}(x)-\mathfrak{B}(s_{0}),\,x-s_{0}\big\rangle_{X}=c\Big\},$
for $c\in\mathbb{R}$.
Proof.
The proof can be found in the Appendix 8.1. ∎
We now are in the position to provide a stability result for the primal-dual dynamics (4.1).
Application to the primal-dual gradient dynamics. As we consider the unconstrained case, i.e. if $F=U$, then the row and column of (4.1) corresponding to the Lagrange multiplier $\mu$ do not occur. Precisely, define
$M_{\mathrm{opt}}^{0}:\begin{bmatrix}\left[\begin{smallmatrix}P\\ U\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}P\\ \mathbb{R}^{n}\end{smallmatrix}\right]\end{bmatrix}\supset\begin{bmatrix}D(\mathcal{C})\\ D(\mathcal{C}^{*})\end{bmatrix}\rightarrow\begin{bmatrix}\left[\begin{smallmatrix}P\\ U\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}P\\ \mathbb{R}^{n}\end{smallmatrix}\right]\end{bmatrix}\eqqcolon\mathcal{W}_{0}$
via
$\underbrace{\begin{bmatrix}\left[\begin{smallmatrix}x\\ u\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}\lambda\\ \lambda_{0}\end{smallmatrix}\right]\end{bmatrix}}_{\eqqcolon w_{0}}\mapsto\underbrace{\begin{bmatrix}\nabla J(x,u)\\ 0\end{bmatrix}}_{\eqqcolon\mathcal{M}_{1}^{0}(w_{0})}+\underbrace{\begin{bmatrix}0&\mathcal{C}^{*}\\ -\mathcal{C}&0\end{bmatrix}}_{\eqqcolon\mathcal{M}_{2}^{0}}\begin{bmatrix}\left[\begin{smallmatrix}x\\ u\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}\lambda\\ \lambda_{0}\end{smallmatrix}\right]\end{bmatrix}$.
Hence, the optimizer dynamics (4.1) reduces to
$\frac{\mathrm{d}}{\mathrm{d}t}w_{0}(t)+M_{\mathrm{opt}}^{0}(w_{0}(t))=\begin{bmatrix}0\\ -I\end{bmatrix}u_{0}$ (4.9)
with input $u_{0}(t)$ corresponding to $\left[\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right]$ and output
$y_{0}(t)=\begin{bmatrix}0&-I\end{bmatrix}w_{0}(t).$
It is clear from the proof of Theorem 3.1 that there exists a unique steady state
$(w_{0}^{\star},u_{0}^{\star})\coloneqq((x^{\star},u^{\star},\lambda,\lambda_{0}),(f,x_{0}))$ (4.10)
of (4.9). The choice of constant input $u_{0}(t)\equiv\left[\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right]$ leads to an autonomous maximally monotone pHs governed by the operator
$\widetilde{M_{\mathrm{opt}}^{0}}(w_{0})=M_{\mathrm{opt}}^{0}(w_{0})+\begin{bmatrix}0\\ \left[\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right]\end{bmatrix}$.
It is immediate from the definition of (maximal) monotonicity that the operator $\widetilde{M_{\mathrm{opt}}^{0}}$ is again maximally monotone.
In what follows, we study the asymptotic behavior of the semigroup associated with the dynamics
$\frac{\mathrm{d}}{\mathrm{d}t}w_{0}(t)+\widetilde{M_{\mathrm{opt}}^{0}}(w_{0}(t))=0.$ (4.11)
If the cost functional $J$ is strictly convex, then the asymptotic convergence of (4.11) towards the optimal point as in Theorem 3.1 was shown in [10, Cor. 3.5]. We now show a similar result by dropping the assumption of strictly convex costs and instead impose a controllability and observability-type assumption on the optimal control problem (3.1).
Assumption 4.5.
Let $(A,B)$ be controllable in the sense of Kalman. Moreover, assume that whenever a solution $x$ of $\dot{x}=Ax+Bu$ fulfills
$\langle\nabla_{x}J(x(t))-\nabla_{x}J(x^{\star}),x(t)-x^{\star}\rangle_{P}=0,$ (4.12)
for $x^{\star}$ as in (4.10) and for all $t\in[0,s]$ where $s>0$ is arbitrary but fixed, then $x(t)=x^{\star}$ for all $t\in[0,s]$.
Remark 4.6.
We briefly comment on the observability notion introduced in the previous Assumption 4.5. For quadratic cost functionals $\ell_{x}(x)=\tfrac{1}{2}\|Cx\|^{2}$ with $C\in\mathbb{R}^{p\times n}$ for $p\in\mathbb{N}$, (4.12) implies that the output satisfies $C(x(t)-x^{*}(t))=0$ for all $t\in[0,s]$. Hence Assumption 4.5 implies that $x(t)=x^{*}(t)$, as usually in observability.
The following result ensures asymptotic stability of the primal-dual gradient dnyamics.
Theorem 4.7.
Let Assumption 4.5 be fulfilled. Then, for all $\varepsilon>0$ and initial values $w_{0}(0)\in Z=B_{\varepsilon}^{\mathcal{W}_{0}}(w_{0}^{\star})$ the solution $w_{0}$ of (4.9) satisfies
$w_{0}(t)\rightharpoonup(x^{\star},u^{\star},0,0)\quad\text{ as $t\to\infty$.}$
Proof.
We successively verify the assumptions of Theorem 4.2. It is well-known that every closed ball $B_{\varepsilon}^{\mathcal{V}_{0}}(w_{0}^{\star})$ in a reflexive Banach space is weakly sequentially compact . To prove the positive invariance of $Z$ with respect to (4.11), let $w_{0}\in Z$. Note that since $(w_{0}^{\star},u_{0}^{\star})$ is a steady state of (4.9) we have that
$w_{0}^{\star}\in\operatorname{zer}\widetilde{M_{\mathrm{opt}}^{0}}=\operatorname{Fix}T_{\widetilde{M_{\mathrm{opt}}^{0}}}$
by [25, Lem. A.9]. Since $(T_{\widetilde{M_{\mathrm{opt}}^{0}}}(t))_{t\geq 0}$ is a semigroup of nonexpansive mappings, we obtain
[…]
for all $t\geq 0$ and $w\in Z$ which shows the positive invariance. It is obvious that $V(\cdot)=\frac{1}{2}||\cdot-w_{0}^{\star}||^{2}_{\mathcal{W}_{0}}$ is weakly lower semi-continuous and since $(T_{\widetilde{M_{\mathrm{opt}}^{0}}}(t))_{t\geq 0}$ defines a family of nonexpansive mappings and $w_{0}^{\star}\in\operatorname{Fix}T_{\widetilde{M_{\mathrm{opt}}^{0}}}$, we immediately get
$\dot{V}_{T_{\widetilde{M_{\mathrm{opt}}^{0}}}}(w)$
[…]
or all $w\in\mathcal{W}_{0}$ and in particular for all $w\in Z$. Then, we have that
$S_{0}=Z\cap\dot{V}_{T_{\widetilde{M_{\mathrm{opt}}^{0}}}}^{-1}(\{0\})$
and we use Theorem 4.4 to conclude
$S_{0}$ $=\left\{w\in Z\,\middle|\,\left\langle\begin{bmatrix}\nabla J\left(\begin{smallmatrix}x\\ u\end{smallmatrix}\right)-\nabla J\left(\begin{smallmatrix}x^{\star}\\ u^{\star}\end{smallmatrix}\right)\\ \left[\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right]-\left[\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right]\end{bmatrix},w-w_{0}^{\star}\right\rangle_{\mathcal{W}_{0}}\right\}$
$=\left\{w\in Z\,\middle|\,\left\langle\nabla J\left(\begin{smallmatrix}x\\ u\end{smallmatrix}\right)-\nabla J\left(\begin{smallmatrix}x^{\star}\\ u^{\star}\end{smallmatrix}\right),\left(\begin{smallmatrix}x\\ u\end{smallmatrix}\right)-\left(\begin{smallmatrix}x^{\star}\\ u^{\star}\end{smallmatrix}\right)\right\rangle\right\}.$
The strict convexity of $l_{u}$ as in (3.2) show that an element $w=(\left(\begin{smallmatrix}x\\ u\end{smallmatrix}\right),\left(\begin{smallmatrix}\lambda\\ \lambda_{0}\end{smallmatrix}\right))$ belongs to $S_{0}$ if and only if
$\langle\nabla_{x}J(x)-\nabla_{x}J(x^{\star}),x-x^{\star}\rangle_{P}=0,\quad\text{and}\quad u=u^{\star}.$
Now, let $w_{0}(t)\in\mathcal{W}_{0}$ be a solution of (4.11) on a finite time interval $[0,s]$ belonging identically to $S_{0}$. Then, the line in (4.11) corresponding to $u$ reads
$0=\tfrac{\mathrm{d}}{\mathrm{d}t}u^{\star}=\tfrac{\mathrm{d}}{\mathrm{d}t}u(t)=-\nabla_{u}J(u^{\star})+B^{\top}\lambda(t).$
As $\nabla_{u}J(u^{\star})$ is independent of $t$ and as controllability of the system $(A,B)$ implies observability of the dual system, we have $\lambda(t)\equiv c$ for some constant $c\in\mathbb{R}^{n}$ on the finite time interval $[0,s]$. But $(\lambda,\lambda_{0})\in D(\mathcal{C}^{*})$, i.e., $\lambda(t)\equiv\lambda(t_{f})=0$ and $\lambda_{0}=\lambda(0)=0$. Consequently, the dynamics (4.11) reduces to
$\tfrac{\mathrm{d}}{\mathrm{d}t}x(t)=-\nabla_{x}J(x(t)),\qquad\mathcal{C}\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}=\begin{bmatrix}f\\ x_{0}\end{bmatrix}$
on $[0,s]$. The equation implies then that a solution of $\dot{x}=Ax+Bu$ satisfies (4.12) on $[0,s]$. Hence, $x(t)=x^{\star}$ and the largest invariant set in $S_{0}$ is the singleton
$s_{0}=(x^{\star},u^{\star},0,0).$
Then, the result follows by Theorem 4.2. ∎
4.2. Stability of constrained optimizer dynamics
In case of inequality constraints, the LaSalle-type invariance principle of Theorem 4.2 is not readily applicable to the set-valued evolution equation (4.1). Assuming strict convexity of the cost we thus provide an alternative proof to establish exponential convergence of the optimizer dynamics (4.1) towards the optimal solution $(x^{\star},u^{\star})$ of problem (3.1).
Theorem 4.8.
Let $J$ as in (3.2) be strictly convex and define $u_{\mathrm{opt}}(t)\equiv\left[\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right]$. Moreover, let $\left(\begin{smallmatrix}x^{\star}\\ u^{\star}\end{smallmatrix}\right)$ be the unique optimal solution of (3.1). Then, the $\left(\begin{smallmatrix}x\\ u\end{smallmatrix}\right)$-component of the strong solution of the autonomous dynamics (4.1) satisfies
$\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}\to\begin{bmatrix}x^{\star}\\ u^{\star}\end{bmatrix}$ (4.13)
exponentially as $t\to\infty$.
Proof.
Since $M_{\mathrm{opt}}$ is maximally monotone and $\mathcal{B}u_{\mathrm{opt}}(t)\equiv\mathcal{B}\left[\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right]\in L^{1}([0,\infty),\mathcal{W})$ it follows that the solution $w\in W^{1,\infty}([0,\infty),\mathcal{W})$ of (4.1) is unique, see the discussion after Definition 2.1. Recall that $(w^{\star},u_{o}^{\star})$ as in (4.2) defines a unique steady state of (4.1). Define the shifted state variable
$\overline{w}\coloneqq w-w^{\star}$
and denote the principal section from (2.5) applied to the maximally monotone operators $M_{\mathrm{opt}}$ and $M_{\mathrm{opt}}-\left[\begin{smallmatrix}\nabla J\\ 0\end{smallmatrix}\right]$ by $M_{\mathrm{opt},0}$ and $M_{\mathrm{opt},J,0}$, respectively. Clearly,
$M_{\mathrm{opt}}=M_{\mathrm{opt}}-\begin{bmatrix}\nabla J\\ 0\end{bmatrix}+\begin{bmatrix}\nabla J\\ 0\end{bmatrix}$
and thus,
$M_{\mathrm{opt},0}=M_{\mathrm{opt},J,0}+\begin{bmatrix}\nabla J\\ 0\end{bmatrix}$
is monotone. For strong solutions $w$, we have
$\tfrac{1}{2}\tfrac{\mathrm{d}}{\mathrm{d}t}\|\overline{w}(t)\|^{2}$ $=\langle-M_{\mathrm{opt},0}(w(t))+\mathcal{B}u_{o}^{\star},\overline{w}(t)\rangle$
$=-\langle M_{\mathrm{opt},0}(w(t))-M_{\mathrm{opt},0}(w^{\star}),\overline{w}(t)\rangle$
$=-\langle M_{\mathrm{opt},J,0}(w(t))-M_{\mathrm{opt},J,0}(w^{\star}),\overline{w}(t)\rangle$
$\hphantom{=}-\left\langle\nabla J\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}-\nabla J\begin{bmatrix}x^{\star}\\ u^{\star}\end{bmatrix},\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}-\begin{bmatrix}x^{\star}\\ u^{\star}\end{bmatrix}\right\rangle$
$\leq-\beta\left\|\begin{bmatrix}x(t)\\ u(t)\end{bmatrix}-\begin{bmatrix}x^{\star}\\ u^{\star}\end{bmatrix}\right\|^{2}\leq 0$
for some $\beta>0$ and almost all $t\geq 0$, by strict convexity of $J$. Hence, the norm is monotonically non-increasing. Furthermore, integration results in
$\left\|\begin{bmatrix}\overline{x}(t)\\ \overline{u}(t)\end{bmatrix}\right\|^{2}-\left\|\overline{w}(0)\right\|^{2}$ $\leq\left\|\overline{w}(t)\right\|^{2}-\left\|\overline{w}(0)\right\|^{2}$
$\leq-2\int_{0}^{t}\beta\left\|\begin{bmatrix}\overline{x}(\tau)\\ \overline{u}(\tau)\end{bmatrix}\right\|^{2}\mathrm{d}\tau,$
which holds for all $t\geq 0$. Applying the Gronwall Lemma from [12, Lemma 4.1] leads to
$\left\|\begin{bmatrix}\overline{x}(t)\\ \overline{u}(t)\end{bmatrix}\right\|\leq\left\|\begin{bmatrix}\overline{x}(0)\\ \overline{u}(0)\end{bmatrix}\right\|e^{-\beta t},\quad\forall\,t\geq 0.$ (4.14)
This implies that the claim. ∎
5. Closed-loop: Suboptimal MPC-type controller
In this section, we introduce a control-by-interconnection scheme that stabilizes the maximally monotone nonlinear pH system
$\dot{x}_{p}(t)$ $=-M_{p}(x_{p}(t))+B_{p}u_{p}(t)$ (5.1)
$y_{p}(t)$ $=B_{p}^{\top}x_{p}(t),$ (5.2)
where $B_{p}\in\mathbb{R}^{n\times m}$ and $M_{p}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a monotone operator. To this end, as illustrated in Figure 2, the above system is coupled with the optimizer dynamics (4.1) in a structure preserving manner. To be precise, we first consider the case where no inequality constraints in the optimal control problem (3.1) appear. The interconnection law is inspired by an MPC feedback, where an initial part of the optimal solution $u^{*}\in U$ is fed back into the plant. Afterwards, we include inequality constraints into our setting.
5.1. The case without control constraints.
Our aim is to interconnect the plant system (5.1) and the optimizer dynamics (4.1) in a structure preserving manner. To this end, the cost functional corresponding to the control is chosen as
$\ell_{u}(u)=\tfrac{\alpha}{2}\|u\|^{2}_{U}$
for some $\alpha>0$. Due to the second line of the optimality system (3.6) and the quadratic control costs, the optimal control $u^{\star}$ satisfies
$\nabla_{u}J(u^{\star})-B^{\top}\lambda=\alpha u^{\star}-B^{\top}\lambda=0$
for some $\lambda$ with $(\lambda,\lambda(0))\in D(\mathcal{C}^{*})$. Thus, in view of Figure 1 the MPC feedback satisfies
$u^{\star}(0)=\tfrac{1}{\alpha}B^{\top}\lambda(0)=\tfrac{1}{\alpha}B^{\top}\lambda_{0}.$
Motivated by this correspondence in nominal MPC we define the suboptimal MPC interconnection law
$\begin{bmatrix}u_{p}\\ u_{o}^{1}\\ u_{o}^{2}\end{bmatrix}=\tfrac{1}{\alpha}\begin{bmatrix}0&0&-B^{\top}\\ 0&0&0\\ B&0&0\end{bmatrix}\begin{bmatrix}y_{p}\\ y_{o}^{1}\\ y_{o}^{2}\end{bmatrix}$ (5.3)
with a splitting $u_{o}=\left[\begin{smallmatrix}u_{o}^{1}\\ u_{o}^{2}\end{smallmatrix}\right]$ corresponding to $\left[\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right]$ and $y_{o}=\left[\begin{smallmatrix}y_{o}^{1}\\ y_{o}^{2}\end{smallmatrix}\right]$ corresponding to $\left[\begin{smallmatrix}-\lambda\\ -\lambda_{0}\end{smallmatrix}\right]$. In order to introduce the coupled system, we define
$\mathcal{V}_{0}\coloneqq\mathbb{R}^{n}\times P\times U\times P\times\mathbb{R}^{n}$ (5.4)
and $N_{0}:\mathcal{V}_{0}\to\mathcal{V}_{0}$ via
$N_{0}\begin{bmatrix}x_{p}\\ \left[\begin{smallmatrix}x\\ u\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}\lambda\\ \lambda_{0}\end{smallmatrix}\right]\end{bmatrix}=\begin{bmatrix}M_{p}(x_{p})\\ \nabla J(x,u)\\ 0\end{bmatrix},\qquad D(N_{0})=\mathcal{V}_{0}.$ (5.5)
Correspondingly, $K:\mathcal{V}_{0}\supset D(K_{0})\to\mathcal{V}_{0}$ is given by
[…]
and
$D(K_{0})=\mathbb{R}^{n}\times D(\mathcal{C})\times D(\mathcal{C}^{*}).$
In this case, we let $M_{0}\coloneqq N_{0}+K_{0}$ and the interconnection (5.3) leads to the autonomous maximally monotone pH system
$\frac{\mathrm{d}}{\mathrm{d}t}v_{0}(t)+M_{0}(v_{0}(t))=0.$ (5.6)
In what follows, we investigate the stability behavior of the coupled system (5.6). Our aim is to use the LaSalle-type weak invariance principle from Theorem 4.2. To this end, we consider $V:\mathcal{V}_{0}\to\mathbb{R}_{+}$ with
$V(v_{0})=\tfrac{1}{2}\|v_{0}\|_{\mathcal{V}_{0}}^{2}$
which clearly defines a Lyapunov functional for (5.6) in the sense of the LaSalle principle of Subsection 4.1.
Lemma 5.1.
Let $M_{p}(0)=0$ and $J$ be strictly convex. Furthermore, let $Z\subset\mathcal{V}_{0}$. Then, the level set $S_{0}$ defined in (4.3) satisfies
$S_{0}=\left\{v_{0}\in Z\,\middle|\,\langle M_{p}(x_{p}),x_{p}\rangle_{\mathbb{R}^{n}}=0,\left[\begin{smallmatrix}x\\ u\end{smallmatrix}\right]=0\right\}.$
Proof.
Observe that
$S_{0}=\dot{V}_{M_{0}}^{-1}(\{0\})\cap Z$
such that Theorem 4.3 implies
$S_{0}=\left\{v_{0}\in Z\,\middle|\,\langle N_{0}(v_{0}),v_{0}\rangle_{\mathcal{V}_{0}}=0\right\}.$
The definition of $N_{0}$ (see (5.5)) thus yields that $v\in S_{0}$ if and only if
$0=\langle M_{p}(x_{p}),x_{p}\rangle_{\mathbb{R}^{n}}+\left\langle\nabla J\begin{bmatrix}x\\ u\end{bmatrix},\begin{bmatrix}x\\ u\end{bmatrix}\right\rangle_{P\times U}.$
However, due to strict convexity of $J$,
$\left\langle\nabla J\begin{bmatrix}x\\ u\end{bmatrix},\begin{bmatrix}x\\ u\end{bmatrix}\right\rangle_{P\times U}\geq c_{0}\left\|\begin{bmatrix}x\\ u\end{bmatrix}\right\|^{2}_{P\times U}$
and thus $v\in S_{0}$ implies that
$0$ $\geq c_{0}\left\|\begin{bmatrix}x\\ u\end{bmatrix}\right\|^{2}_{P\times U}+\langle M_{p}(x_{p}),x_{p}\rangle_{\mathbb{R}^{n}}$
$=c_{0}\left\|\begin{bmatrix}x\\ u\end{bmatrix}\right\|^{2}_{P\times U}+\langle M_{p}(x_{p})-M_{p}(0),x_{p}-0\rangle_{\mathbb{R}^{n}}.$
As $M_{p}$ is monotone, the second term is nonnegative which implies that $\langle M_{p}(x_{p}),x_{p}\rangle_{\mathbb{R}^{n}}=0$ and $(x,u)=0$. ∎
Theorem 5.2.
Assume that the plant dynamics (5.1) is observable, i.e. if a state trajectory $x_{p}:[0,s]\to\mathbb{R}^{n}$ satisfies $B_{p}^{\top}x_{p}(t)=0$ for $t\in[0,s]$, then $x_{p}(t)=0$ on $[0,s]$. Moreover, let $M_{p}(0)=0$ and $J$ be strictly convex. Then for all $\varepsilon>0$ and initial values $v_{0}(0)\in Z=B_{\varepsilon}^{\mathcal{V}_{0}}(0)$ the solution $v_{0}(t)$ of (5.6) satisfies
$v_{0}(t)\rightharpoonup 0$
as $t\to\infty$. In particular, $x_{p}(t)\rightarrow 0$ as $t\rightarrow 0$.
Proof.
We prove the stability result by applying the invariance principle 4.2. To this end, we first show that the largest invariant set in $S_{0}$ only contains the origin.
Let $v_{0}(t)\in\mathcal{V}_{0}$ be a solution of (5.6) on a finite time interval $[0,s]$ belonging identically to $S_{0}$. Thus,
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}x_{p}(t)\\ \left[\begin{smallmatrix}0\\ 0\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}\lambda(t)\\ \lambda_{0}(t)\end{smallmatrix}\right]\end{bmatrix}=\begin{bmatrix}-M_{p}(x_{p}(t))+\tfrac{1}{\alpha}B_{p}B^{\top}\lambda_{0}(t)\\ -\mathcal{C}^{*}\left[\begin{smallmatrix}\lambda(t)\\ \lambda_{0}(t)\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}0\\ -\tfrac{1}{\alpha}BB_{p}^{\top}x_{p}(t)\end{smallmatrix}\right]\end{bmatrix}$
where $x_{p}(t)$ satisfies $\langle M_{p}(x_{p}(t)),x_{p}(t)\rangle_{\mathbb{R}^{n}}=0$ for all $t\in[0,s]$. The second line of the above equation implies that
$\left[\begin{smallmatrix}\lambda(t)\\ \lambda_{0}(t)\end{smallmatrix}\right]\in\ker\mathcal{C}^{*}=\{0\}$
for all $t\in[0,s]$ where we used [10, Lem. 3.2]. By the third line, we further have that $0=\tfrac{1}{\alpha}BB_{p}^{\top}x_{p}(t)$ for all $t\in[0,s]$ which yields $x_{p}(t)\in\ker B_{p}^{\top}$ by injectivity of $B$. Consequently, the observability assumption implies that $x_{p}(t)=0$ on $[0,s]$ implying that the largest invariant set in $S_{0}$ is $\{0_{\mathcal{V}_{0}}\}$.
To show the asymptotic stability, we successively verify the assumptions of Theorem 4.2. It is well-known that every closed ball $B_{\varepsilon}^{\mathcal{V}_{0}}(0)$ is weakly sequentially compact in a reflexive Banach space. To prove the positive invariance of $Z$ with respect to (5.6), let $v_{0}\in Z$. Note that $N_{0}(0)=0$ by assumption. Hence, $x(t)\equiv 0$ is a solution of (5.6) on $0,\infty)$. The uniqueness of mild solutions then gives $T_{M_{0}}(t)(0)=0$ for all $t\geq 0$. Since $T_{M_{0}}(t))_{t\geq 0}$ is semigroup of nonexpansive mappings, we have
[…]
for all $t\geq 0$ which shows the positive invariance. It is obvious that $V(\cdot)=\frac{1}{2}||\cdot||^{2}_{\mathcal{V}_{0}}$ is weakly lower semi-continuous and since $(T_{M_{0}}(t))_{t\geq 0}$ is a family of nonexpansive mappings, we immediately get
[…]
for all $v_{0}\in\mathcal{V}_{0}$ and in particular for all $v_{0}\in Z$. Consequently, the weak invariance principle 4.2 yields that for all $v_{0}\in Z$ we have
$T_{M_{0}}(t)(v_{0})\rightharpoonup 0$
as $t\to\infty$. The weak convergence and the strong convergence are equivalent in finite-dimensional Hilbert spaces and thus, $x_{p}(t)\to 0$ strongly as $t\to\infty$. ∎
The following result allows to drop the strict convexity assumption on the $J$ with respect to $x$.
Theorem 5.3.
Let the assumptions of Theorem 5.2 hold and replace the strict convexity assumption of $J$ with the controllability and detectability conditions on the OCP given in Assumption 4.5. Moreover, let $J$ satisfy $\nabla_{x}J(0)=0$. Then, for all $\varepsilon>0$ and initial values $v_{0}(0)\in Z=B_{\varepsilon}^{\mathcal{V}_{0}}(0)$ the solution $v_{0}(t)$ of (5.6) satisfies
$v_{0}(t)\rightharpoonup 0$
as $t\to\infty$. In particular, $x_{p}(t)\rightarrow 0$ as $t\rightarrow 0$.
Proof.
Following the proof of Theorem 5.2 it is clear that the assumptions of Theorem 4.2 are fulfilled. It remains to show that the largest invariant set in $S_{0}$ is the singleton $\{0_{\mathcal{V}_{0}}\}$. To this end, let $v_{0}(t)\in\mathcal{V}_{0}$ be a solution of (5.6) on a finite time interval $[0,s]$ belonging identically to $S_{0}$. Precisely,
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}x_{p}(t)\\ \left[\begin{smallmatrix}x(t)\\ 0\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}\lambda(t)\\ \lambda_{0}(t)\end{smallmatrix}\right]\end{bmatrix}=\begin{bmatrix}-M_{p}(x_{p}(t))+\tfrac{1}{\alpha}B_{p}B^{\top}\lambda_{0}(t)\\ \left[\begin{smallmatrix}-\nabla_{x}J(x(t),0)\\ 0\end{smallmatrix}\right]-\mathcal{C}^{*}\left[\begin{smallmatrix}\lambda(t)\\ \lambda_{0}(t)\end{smallmatrix}\right]\\ \left[\begin{smallmatrix}0\\ -\tfrac{1}{\alpha}BB_{p}^{\top}x_{p}(t)\end{smallmatrix}\right]+\mathcal{C}\left[\begin{smallmatrix}x(t)\\ 0\end{smallmatrix}\right]\end{bmatrix}$ (5.7)
where $x_{p}(t),x(t)$ satisfy
$\langle M_{p}(x_{p}(t)),x_{p}(t)\rangle_{\mathbb{R}^{n}}=\langle\nabla_{x}J(x(t)),x(t)\rangle_{P}=0$
for all $t\in[0,s]$. Here, monotonicity of $M_{p}$ and $\nabla_{x}J$, together with the fact that both operators vanish at the origin, implies that each term must vanish individually. Note that $u\equiv 0$ on $[0,s]$ follows immediately from strict convexity of $u\mapsto\tfrac{\alpha}{2}\|u\|_{U}^{2}$. It is a well-known fact in systems and control theory that the controllability of $(A,B)$ is equivalent to the observability of the dual system
$\tfrac{\mathrm{d}}{\mathrm{d}\tau}\lambda(\tau)$ $=-A^{\top}\lambda(\tau),\qquad\lambda(0)=\lambda_{0}$
$v(\tau)$ $=-B^{\top}\lambda(\tau).$
The third line of (5.7) implies that $0=B^{\top}\lambda(t)$ for all $t\in[0,s]$, which yields that $\lambda\equiv 0$ on $[0,s]$ by the assumed observability of the adjoint system. Since $\left(\begin{smallmatrix}\lambda\\ \lambda_{0}\end{smallmatrix}\right)\in D(\mathcal{C}^{*})$ we conclude $\lambda_{0}=\lambda(0)=0$. As a consequence, the system dynamics (5.7) reduces to the constrained dynamics
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}x_{p}(t)\\ x(t)\end{bmatrix}$ $=\begin{bmatrix}-M_{p}(x_{p}(t))\\ -\nabla_{x}J(x(t))\end{bmatrix}$,
$\begin{bmatrix}0\\ \tfrac{1}{\alpha}BB_{p}^{\top}\end{bmatrix}x_{p}(t)$ $=\begin{bmatrix}\tfrac{\mathrm{d}}{\mathrm{d}\tau}x(t)-Ax(t)\\ x(t)(0)\end{bmatrix}$.
The constraint implies
$x(t)=e^{A\cdot}x(t)(0)=\tfrac{1}{\alpha}e^{A\cdot}BB_{p}^{\top}x_{p}(t)$
since $x(t)$ solves the ordinary differential equation
$\tfrac{\mathrm{d}}{\mathrm{d}\tau}x(\tau)=Ax(\tau),\qquad x(0)=\tfrac{1}{\alpha}BB_{p}^{\top}x_{p}(t)$
for each time $t\in[0,s]$. Consequently,
$0$ $=\langle x(t),\nabla_{x}J(x(t))\rangle_{P}$
$=\langle\tfrac{1}{\alpha}e^{A\cdot}BB_{p}^{\top}x_{p}(t),\nabla_{x}J(\tfrac{1}{\alpha}e^{A\cdot}BB_{p}^{\top}x_{p}(t),0\rangle_{P}$
and thus $\tfrac{1}{\alpha}e^{A\cdot}BB_{p}^{\top}x_{p}(t)=0$ for all $t\in[0,s]$, where we utilized the supposed observability. With injectivity of $B$ we conclude that $B_{p}^{\top}x_{p}(t)=0$ for all $t\in[0,s]$ and hence, $x_{p}(t)=0$ due to the observability of the plant. The remaining part of the proof is analogous to the proof of Theorem 5.2. ∎
The observability notion suggested in Assumption 4.5 can be specified in the case of quadratic costs.
Remark 5.4 (Quadratic costs).
Assume that the cost functional $J:P\times U\rightarrow\mathbb{R}$ is given by
[…]
$\nabla J(x,u)=\begin{bmatrix}C^{\top}Cx\\ \alpha u\end{bmatrix}$.
It is easily seen that the observability assumption (4.12) reduces to the standard observability of the pair $(A,C)$, where $A$ is seen as a bounded multiplication operator on $P$. Precisely, in this case Assumption 4.5 is equivalent to the following: for any solution $x$ of the state-output system
$\tfrac{\mathrm{d}}{\mathrm{d}t}x(t)$ $=Ax(t),\qquad x(0)=x_{0}$
$y(t)$ $=Cx(t)$
on $[0,s]$, the equality
$Cx(t)=0$
implies that $x\equiv 0$ in $[0,s]$.
5.2. Inequality constrained problems
Throughout this subsection, we assume that the gradient of the cost function $J$ with respect to $x$ vanishes at the origin. If $F$ is given as in (3.4), then by (adjoint), the optimal control $u^{\star}$ satisfies
$\alpha u^{\star}-B^{\top}\lambda^{\star}+\mu=0$
for some $\mu\in\partial i_{F}(u^{\star})$. We conclude
$\tfrac{1}{\alpha}B^{\top}\lambda-u^{\star}\in\tfrac{1}{\alpha}\partial i_{F}(u^{\star})=\partial\tfrac{1}{\alpha}i_{F}(u^{\star})$
and thus,
$u^{\star}=(I+\partial\tfrac{1}{\alpha}i_{F})^{-1}(\tfrac{1}{\alpha}B^{\top}\lambda)\eqqcolon\operatorname{prox}_{\tfrac{1}{\alpha}i_{F}}(\tfrac{1}{\alpha}B^{\top}\lambda)$
by [2, Prop. 16.44]. But since the scaling $\tfrac{1}{\alpha}$ does not affect the indicator function $i_{F}$ we have the formula
$u^{\star}=\operatorname{prox}_{i_{F}}(\tfrac{1}{\alpha}B^{\top}\lambda)=P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda)$ (5.9)
where $P_{F}$ denotes the projection onto the set $F$, [2, Ex. 12.25]. In MPC, the initial part
$u^{\star}(0)=\operatorname{prox}_{i_{F}}(\tfrac{1}{\alpha}B^{\top}\lambda(0))=\operatorname{prox}_{i_{F}}(\tfrac{1}{\alpha}B^{\top}\lambda_{0})$
of the optimal control is fed back into the plant. Our approach will be a combination of this MPC feedback and a so-called saturated output control
$u=-\operatorname{prox}_{i_{F}}(B_{p}^{\top}x_{p})$
that is common in control of pH systems, see [30]. Hence, in the spirit of suboptimal MPC and saturating output control, we couple the systems (5.1) and (4.1) via the structure preserving interconnection
$\begin{bmatrix}u_{p}\\ u_{0}\end{bmatrix}=\begin{bmatrix}-c&0\\ 0&c\end{bmatrix}\begin{bmatrix}\mathcal{F}_{1}^{*}\mathcal{E}(\mathcal{F}_{1}y_{p})-\mathcal{F}_{1}^{*}\mathcal{E}(\mathcal{F}_{2}y_{o})\\ \mathcal{F}_{2}^{*}\mathcal{E}(\mathcal{F}_{1}y_{p})-\mathcal{F}_{2}^{*}\mathcal{E}(\mathcal{F}_{2}y_{o})\end{bmatrix}$, (5.10)
with the partitioning $u_{o}=\left[\begin{smallmatrix}u_{o}^{1}\\ u_{o}^{2}\end{smallmatrix}\right]$ corresponding to $\left[\begin{smallmatrix}f\\ x_{0}\end{smallmatrix}\right]$, $y=\left[\begin{smallmatrix}y_{1}\\ y_{2}\end{smallmatrix}\right]=-\left[\begin{smallmatrix}\lambda\\ \lambda_{0}\end{smallmatrix}\right]$ and where $c=\tfrac{1}{2}$, $\mathcal{E},\mathcal{F}_{1}:U\to U,\mathcal{F}_{2}:P\times\mathbb{R}^{n}\to U$ are defined by
$\mathcal{E}=\operatorname{prox}_{i_{F}},\quad\mathcal{F}_{1}=I,\quad\mathcal{F}_{2}=\tfrac{1}{\alpha}\begin{bmatrix}0&-B^{\top}\end{bmatrix}$.
Inserting the definition of the outputs of the plant (5.1) and optimizer dynamics (4.1) yields
$\begin{bmatrix}u_{p}\\ u_{o}^{1}\\ u_{o}^{2}\end{bmatrix}=\begin{bmatrix}\tfrac{1}{2}P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda_{0})-\tfrac{1}{2}P_{F}(B_{p}^{\top}x_{p})\\ 0\\ \tfrac{1}{2\alpha}BP_{F}(\tfrac{1}{\alpha}B^{\top}\lambda_{0})-\tfrac{1}{2\alpha}BP_{F}(B_{p}^{\top}x_{p})\end{bmatrix}$. (5.11)
Proposition 5.5.
The control $u_{p}$ as in (5.11) is feasible, i.e. $u_{p}(t)\in F$ for all $t>0$.
Proof.
By definition of the projection operator $\operatorname{prox}_{i_{F}}=P_{F}$ we have $P_{F}(x)\in F$ for all $x\in U$. The convexity of $F$ together with the fact F=FF=-F yields
$\vartheta P_{F}(x)-(1-\vartheta)P_{F}(y)\in F$
for all $\vartheta\in[0,1]$ and $x,y\in U$. The choice $\vartheta=\tfrac{1}{2}$ implies the claim. ∎
In view of (5.9), we introduce the control-reduced coupled optimizer plant dynamics on $\mathcal{V}_{\mathrm{r}}\coloneqq\mathbb{R}^{n}\times X\times X\times\mathbb{R}^{n}$
$\frac{\mathrm{d}}{\mathrm{d}t}v(t)=-M_{\mathrm{r}}(v(t))$ (5.12)
where $v(t)=(x_{p}(t),x(t),(\lambda(t),\lambda_{0}(t)))$ and
$M_{\mathrm{r}}(v)=\begin{bmatrix}M_{p}(x_{p})-\tfrac{1}{2}B_{p}(P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda_{0})-P_{F}(B_{p}^{\top}x_{p}))\\ \nabla_{x}J(x)+(-\tfrac{\mathrm{d}}{\mathrm{d}\tau}-A^{\top})\lambda\\ -(\tfrac{\mathrm{d}}{\mathrm{d}\tau}-A)x+BP_{F}(\tfrac{1}{\alpha}B^{\top}\lambda)\\[3.44444pt] \tfrac{1}{2\alpha}B(P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda_{0})-P_{F}(B_{p}^{\top}x_{p}))-x(0)\end{bmatrix}$
with
$D(M_{\mathrm{r}})=\mathbb{R}^{n}\times H^{1}([0,t_{f}],\mathbb{R}^{n})\times D(\mathcal{C}^{*}).$
By Proposition 2.4, the interconnection (5.11) leads to an autonomous maximally monotone pH system (5.12) on $\mathcal{V}_{\mathrm{r}}$.
Theorem 5.6.
Assume that the plant dynamics (5.1) is observable and $\nabla_{x}J(0)=0$. Then, for all $\varepsilon>0$ and initial values $v_{0}\in Z=B_{\varepsilon}^{\mathcal{V}_{\mathrm{r}}}(0)$ the solution $v_{0}(t)$ of (5.12) satisfies $x_{p}(t)\rightarrow 0$ as $t\rightarrow 0$.
Proof.
Again, the proof relies on Theorem 4.2, and the assumptions are verified as in Theorem 4.7. Thusly, it remains to show that the largest invariant set in
$S_{0}=\left\{v\in\mathcal{V}_{\mathrm{r}}\,\middle|\,\langle M_{\mathrm{r}}(v),v\rangle_{\mathcal{V}_{\mathrm{r}}}=0\right\}$
is the zero solution. Observe that
$\langle M_{\mathrm{r}}(v),v\rangle_{\mathcal{V}_{\mathrm{r}}}$ $=\langle M_{p}(x_{p}),x_{p}\rangle_{\mathbb{R}^{n}}+\langle\nabla_{x}J(x),x\rangle_{X}+\alpha\langle P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda),\tfrac{1}{\alpha}B^{\top}\lambda\rangle_{U}$
$\hphantom{=}+\tfrac{1}{2}\langle P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda_{0})-P_{F}(B_{p}^{\top}x_{p}),\tfrac{1}{\alpha}B^{\top}\lambda_{0}-B_{p}^{\top}x_{p}\rangle_{U}$
$\geq\langle M_{p}(x_{p}),x_{p}\rangle_{\mathbb{R}^{n}}+\langle\nabla_{x}J(x),x\rangle_{X}+\alpha\|P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda)\|^{2}_{U}$
$\hphantom{\geq}+\tfrac{1}{2}\|P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda_{0})-P_{F}(B_{p}^{\top}x_{p})\|_{U}^{2}$
where we used that the projection operator $P_{F}$ is firmly nonexpansive, cf. (2.10). Since each of the above terms is lower bounded by 0 we conclude $(x_{p},x,\lambda,\lambda_{0})\in S_{0}$ if and only if
$\langle M_{p}(x_{p}),x_{p}\rangle_{\mathbb{R}^{n}}=\langle\nabla_{x}J(x),x\rangle_{X}=\|P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda)\|_{U}^{2}=0$
and $\|P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda_{0})-P_{F}(B_{p}^{\top}x_{p})\|^{2}=0$. Let $v(t)$ be a solution of (5.12) that is contained in $S_{0}$. Then,
$P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda(t))\equiv 0\qquad\Longrightarrow\qquad P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda_{0})=P_{F}(\tfrac{1}{\alpha}B^{\top}\lambda(0))=0$
and hence,
$P_{F}(B_{p}^{\top}x_{p}(t))\equiv 0\qquad\xRightarrow{0\in\operatorname{int}F}\qquad B_{p}^{\top}x_{p}(t)=0.$
The observability of the plant yields then $x_{p}(t)\to 0$. ∎
6. Numerical examples
We consider a nonlinear energy conserving system. Precisely, we consider the plant dynamics
[…]
$y(t)$ $=z_{2}(t)$
The governing operator $M_{p}$ is energy conserving in the sense that
$\langle M_{p}(z),z\rangle_{\mathbb{R}^{2}}=0$
for all $z\in\mathbb{R}^{2}$. Consequently, any solution of the uncontrolled system preserves its norm over time. We introduce the optimization problem
[…]
on a finite time horizon $t_{f}>0$ that underlies the proposed control-by-interconnection scheme. Depending on whether we want to incorporate control constraints into our setting, we require the optimal control $u\in L^{2}([0,t_{f}],\mathbb{R})$ to be contained in the set
$F=\left\{u\in L^{2}([0,t_{f}],\mathbb{R})\,\middle|\,\underline{u}\leq u(t)\leq\overline{u}\ \text{a.e. on}\ [0,t_{f}]\right\}$
for some choice $\overline{u}=-\underline{u}\in\mathbb{R}_{+}$. We approach the two evolution equations (5.6) (unconstrained case) and (5.12) (constrained case) numerically. To this end, the differential operators $\mathcal{C}$ and $\mathcal{C}^{*}$ are approximated by the implicit Euler rule and the resulting finite-dimensional system is solved using an explicit Runge-Kutta method with a very fine resolution. The norm of the plant solution over time is shown in Figure 3. The feedback controls that are applied to the plant can be seen in Figure 4. It can be observed that the control in the constrained case is pointwisely contained in the feasible set $F$. The norm of the solution of the full coupled system over time is illustrated in 5. Observe that, due to monotonicity of the governing operators, the map $t\mapsto\|w(t)\|^{2}$ is monotonically nonincreasing. Indeed, the numerical approximation of the fully coupled system is observed to be asymptotically stable, indicating convergence of the Lagrange multipliers $(\lambda,\lambda_{0},\mu)(t)$ to the unique optimal point of the underlying optimality system (3.8).
Figure 3. Norm of the state component corresponding to the plant with initial value $x_{0}=(-0.5,-3)^{\top}$,
Figure 4. Visualization of plant inputs on symmetric logarithmic scale: $u_{\mathrm{unconstr}}=\tfrac{1}{\alpha}B^{\top}\lambda_{0}$ as in (5.3) and saturated input $u_{\mathrm{constr}}$ as in (5.11) with $\overline{u}=-\underline{u}=1$, $\alpha=1.5$
Figure 5. Norm of the coupled optimizer-plant dynamics over time.
7. Conclusion
We proposed an optimization-based control-by-interconnection framework for the stabilization of nonlinear port-Hamiltonian systems. By interpreting a primal–dual gradient dynamics associated with a finite-horizon, possibly control-constrained optimal control problem as a port-Hamiltonian system, we obtained an MPC-type feedback law that naturally preserves the underlying system structure. The resulting closed-loop dynamics arise from a power-preserving interconnection of two port-Hamiltonian systems, namely the plant and the optimization dynamics. Under a suitable observability assumption, we established asymptotic stability of the interconnected system. The effectiveness of the proposed approach was demonstrated through a numerical example. These results provide a systematic bridge between model predictive control and structure-preserving control design for port-Hamiltonian systems.